### v1

@@ 3.2 Optimal experimental design: main result
See Appendix B.1. ∎
Theorem 1 provides an explicit expression of the regret (which does not require researchers to specify $B$). To our knowledge, it is the first result to study adaptation regret for experimental design. Its derivation requires us to find precise binding patterns as a function of $B$, which accounts for its effect on optimizing the oracle solution over the design and estimator.
The adaptation regret balances the variance of a given design relative to the smallest possible variance $\alpha^{\star}$ against the worst-case bias $\beta$ relative to the smallest possible worst-case bias $\beta^{\star}$.
To gain insight into the optimal solution, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), we may expect the minimizer $(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})\in\arg\min\mathcal{R}(\mathcal{E},\Sigma,\gamma)$ to equalize
$\frac{\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})}{\alpha^{\star}}\;=\;\frac{\beta(\mathcal{E}^{\star},\gamma^{\star})}{\beta^{\star}}.$
That is, at the optimum the design equalizes the estimator’s variance $\mathbb{V}(\hat{\tau}_{\gamma})$ (over the design choice) divided by the smallest feasible variance $\alpha^{\star}$ to the worst-case squared bias (equal to $\beta B^{2}$) divided by the smallest achievable worst-case squared bias $\beta^{\star}B^{2}$.
@@
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot), i.e., where the two normalized components are equalized: $\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})/\alpha^{\star}=\beta(\mathcal{E}^{\star},\gamma^{\star})/\beta^{\star}$.
Theorem 1 provides us an immediate and simple to compute solution to the optimal design problem via backward induction. Specifically, we compute the estimator, the optimal variance and the experimental choice as follows:
@@
Optimization proceeds as follows:
  • we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. Importantly, the choice of $\gamma^{\star}$ depends on the class of designs available to researchers ($\mathcal{D}$), because $\mathcal{D}$ determines the oracle values $\alpha^{\star},\beta^{\star}$ in the denominators of the objective function (a larger class of designs improves the oracle solution);
  • the choice of the experimental variance $\Sigma$ (how precise each experiment is under budget constraints) also depends on the bias component $\beta$ because $\beta$ affects how we select the shrinkage weights $\gamma^{\star}$ in the previous step;
  • the optimal design $\mathcal{E}^{\star}$ then accounts for both choices of $\gamma^{\star}$ and $\Sigma^{\star}$.
3.3 Example with two parameters
Consider the two-parameters model of Setting 2 where for simplicity we fix the choice of the variances $\sigma^{2},v^{2}>0$. The smallest variance and worst–case squared bias (per unit $B^{2}$) are
$\alpha^{\star}=\min_{k\in\{1,2\}}\underbrace{\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\}}_{\text{Variance with variance-optimal }\gamma},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\underbrace{\omega_{k}^{2}}_{\text{Bias}^{2}/B^{2}\text{ with bias-optimal }\gamma}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ follows from the fact that the oracle that minimizes the worst-case bias chooses the experiment $k$ with the largest $\omega_{k}$, sets $\gamma_{k}=1$ and incurs a bias proportional to $|\omega_{-k}|$, since no experiment is conducted for k-k.
The variance and worst-case squared bias (divided by $B^{2}$) incurred by an analyst choosing experiment $j$ and $\gamma_{j}$ are, respectively
@@ 3.3 Example with two parameters
By Theorem 1, the worst-case adaptation regret for choosing experiment $\{j\}$ and $\gamma_{j}$, equals
$\mathcal{R}(\{j\},\gamma_{j})=\max\Big\{\underbrace{\alpha(j,\gamma_{j})/\alpha^{\star}}_{\text{Variance/}\alpha^{\star}},\ \underbrace{\beta(j,\gamma_{j})/\beta^{\star}}_{\text{Bias}^{2}/(B^{2}\beta^{\star})}\Big\}.$
Both maximand components are evaluated with the same choice of $\gamma$.
3.3.1 Choosing $\gamma_{j}$ for given design $\{j\}$
The first step is to choose the shrinkage weight $\gamma$, noting that regret is computed relative to an oracle that optimizes both $\gamma$ and the experiment design $\{j\}$. Minimizing $\mathcal{R}$ over $\gamma_{j}$ yields a rule that either equalizes the two components or selects a boundary solution.
Corollary 1.
Consider Setting 2 and let Assumption 1 hold. Then
@@ 3.3.1 Choosing $\gamma_{j}$ for given design $\{j\}$
At the boundaries, if the relative variance $\alpha/\alpha^{\star}$ is always smaller than the relative bias $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance term always dominates the bias term, the optimal choice is the variance–minimizing weight $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$. In most applications, the solution is interior: we choose $\gamma_{j}^{\star}$ strictly between the boundaries to equalize the two contributions, $\frac{\alpha(j,\gamma_{j}^{\star})}{\alpha^{\star}}=\frac{\beta(j,\gamma_{j}^{\star})}{\beta^{\star}},$ so that the variance ratio matches the worst–case squared bias ratio.
The optimal solution depends not only on the features of the experiment design $j$ but also on the features of the other feasible design j-j ($v_{-j}^{2},\sigma_{-j}^{2}$ and $\omega_{-j}$) through $\alpha^{\star}$ and $\beta^{\star}$. This differs from choosing the optimal $\gamma$ for a single fixed design, since the choice of $\gamma$ will then interact with the choice of the class of design the researcher (and oracle) can choose.
Illustration in Figure 2
Figure 2 illustrates three cases for $j=2$. In the top row, we plot $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
Table 1: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$. Each row summarizes each observation 1-4 for $\gamma_{2}^{\star}$.
Figure 2: Illustration for experiment $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
3.3.2 Choosing the design $\{j\}$
After computing $\gamma_{j}^{\star}$, we then optimize over the design index $j$, by choosing
$j^{\star}\in\arg\min_{j\in\{1,2\}}\max\Big\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\ \beta(j,\gamma_{j}^{\star})/\beta^{\star}\Big\}.$ (6)
To build intuition, suppose $\gamma_{j}^{\star}$ is interior for both $j\in\{1,2\}$. Then $\alpha(j,\gamma_{j}^{\star})/\alpha^{\star}=\beta(j,\gamma_{j}^{\star})/\beta^{\star}$, which implies
In this case, optimal $\gamma_{j}^{\star}$ balances variance and worst–case bias. Because $\alpha^{\star}$ and $\beta^{\star}$ do not depend on $j$, $j^{\star}$ minimizes either (and therefore both) criteria $\alpha(j,\gamma_{j}^{\star})$ and $\beta(j,\gamma_{j}^{\star})$.
A second case arises when one experiment (say $j=1$) is bias–dominant (i.e., $\gamma_{1}^{\star}=1$) or variance–dominant (i.e., $\gamma_{1}^{\star}=\sigma_{1}^{2}/(\sigma_{1}^{2}+v_{1}^{2})$), while the other admits an interior solution $\gamma_{2}^{\star}$ for which $\beta(2,\gamma_{2}^{\star})/\beta^{\star}=\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$. In this case, the optimal design selects the experiment that yields the smaller value of the dominating criterion (worst–case bias in the bias–dominant scenario, variance in the variance–dominant scenario).
Although we view these as the leading scenarios, cases at the extreme boundaries can occur. The first is when the relative variance $\alpha/\alpha^{\star}$ or the worst–case bias $\beta/\beta^{\star}$ uniformly dominates the other for both experiments. Then the optimal design minimizes the dominating term; alternatively, if for one experiment $j$ the variance ratio $\alpha/\alpha^{\star}$ is dominant while for the other experiment j-j the worst–case bias ratio $\beta/\beta^{\star}$ is dominant, the minimax regret is the smaller of the two dominating values. These are desiderable properties as the optimal design always prioritizes the dominating term. Table 2 summarizes the cases.
Table 2: Optimal shrinkage and experiment choice across different regimes. The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
Leading cases
Other (boundary) solutions
Illustration in Figure 3
For illustration, Figure 3 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios. The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$. In each column, we keep the same parameterizations as in Figure 2.
Table 3: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 3. Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
Trend regret $j=1$ regret $j=2$ regret $j=1$ regret $j=2$
$\uparrow~v_{2}$ decrease increase decrease increase
$\uparrow~\sigma_{2}$ slow decrease decrease fast increase decrease

### v2

@@ 3.1 Optimal design
Consider Setting 1 and let Assumptions  1, 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$,
@@
See Appendix A.1. ∎
Theorem 1 provides a simple expression of the regret which does not require researchers to specify $B$. The key insight is that the adaptation regret in our context, leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret. This provides a simple and intuitive trade-off between the two.
Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize the variance and bias regret as in Figure 1
$\frac{\alpha}{\alpha^{\star}}\;=\;\frac{\beta}{\beta^{\star}}.$
In practice, this amount to allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
@@
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot).
Theorem 1 provides us an immediate solution to the optimal design problem. We compute the estimator, the optimal variance and the experimental choice via backward induction:
@@
That is, we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. We then choose $\Sigma$, whose choice depends on $\beta$ through the shrinkage weights. We finally choose the set of experiments $\mathcal{E}$.
3.2 Intuition with two parameters
To build further intuition, it is useful to study the two-parameter model.
Setting 2 (Illustration with two-parameter model).
Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed. We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$. If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
How do the variance and bias regret look like in this easier scenario? We can write
$\alpha^{\star}=\min_{k\in\{1,2\}}\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\omega_{k}^{2}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ equals the smallest between the two sensitivity parameters $\omega$. Both depend on the choice of the estimator and on the class of experiments.
We cannot achieve both $\alpha^{\star}$ and $\beta^{\star}$, since these require different choices of $\gamma$. Instead,
@@ 3.2 Intuition with two parameters
The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix C.1).
@@
At the boundaries, if the variance regret $\alpha/\alpha^{\star}$ is always smaller than the bias regret $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance regret always dominates, the optimal choice is the variance-optimal $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
When the solution is an interior solution for $\gamma$, the optimal design minimizes both the bias and variance regret (now equalized), so that
However, cases at the extreme boundaries can also occur. For instance, if the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term. These are desiderable properties as the optimal design always prioritizes the dominating term. Appendix Table 5 provides a complete overview of all possible scenarios. More complex scenarios where researchers may allocate sample sizes to each experiment are possible, see Section 6.
Figure 2: Example for $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Figure 3: Example for $j=2$ (regret). Different colors correspond to the regret for choosing either experiment. Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Example 4 (Numerical illustration).
For an illustrative example, consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is twice as precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$. The optimal $\gamma_{2}^{\star}$ depends not only on features of design $j$ but also on those of the alternative design j-j, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the benchmarks $(\alpha^{\star},\beta^{\star})$. Figure 2 (top row) plots $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. The bottom row reports the corresponding $\gamma_{2}^{\star}$. Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons—when $\omega_{2}$ is small the oracle bias decreases faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ has only on the feasible experiment’s bias. Consequently, both regret and the optimal estimator depend on the oracle’s choice of the experiment. In the interior region, the optimal weight first shifts toward the variance–optimal solution (as the oracle bias declines) and then rises again toward a bias–preferred solution. When we vary the experimental and observational variances, we observe the expected patterns: $\gamma_{2}^{\star}$ moves inversely with experimental variance and in the same direction as observational variance. These patterns translate into implications for the regret of the chosen experiment (Figure 3). For the second arm $j=2$, regret decreases sharply as $\omega_{2}$ increases and then decreases, with further decreases becoming gradual once the bias component is already small. By contrast, for the first arm, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate), but eventually increases with $\omega_{2}$ because the oracle increasingly favors arm $2$, so the opportunity cost of selecting arm $1$ increases. A summary appears in Table 4, with a complete discussion in Appendix C.2. ∎
4 Generalization to moments selection
In many scenarios, we may expect that experiments only identify relevant moments in the data. We now extend the framework to generalized method of moments (GMM) estimators.
Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks all moments (observational and experimental) potentially available for estimation. Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$). For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$). This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.1, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates. For a given covariance matrix $\Sigma$ of the sample moment vector, let $\bar{g}_{\Sigma}$ denote the sample analogue with
$\mathbb{E}[\bar{g}_{\Sigma}]=b,\qquad\mathbb{V}(\bar{g}_{\Sigma})=\Sigma,$
where the vector $b$ captures the biases in the moments. The researcher may decide to access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
Misspecification set.
Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified). We posit bounded misspecification
$\mathcal{B}(B)\;=\;\Big\{b\in\mathbb{R}^{p}:\ b_{j}=0\ \text{for }j\in\mathcal{I},\ \ |b_{j}|\leq B\ \text{for }j\in\mathcal{I}^{c}\Big\}.$ (4)
This restriction states that experimental moments have no bias, whereas observational moments have bias bounded by $B$, potentially unknown to the researcher.
Estimator and design.
Given the empirical moments $\bar{g}_{\Sigma}$, we consider (asymptotically) linear estimators of the form
$\hat{\theta}(W,\Sigma)-\theta\;=\;\Gamma_{\Lambda}(W)\,\bar{g}_{\Sigma},\quad\Gamma_{\Lambda}(W)\equiv-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W$ (5)
where $W$ and $\Sigma$ are chosen by the researcher from a feasible set. The structure of the estimator follows a standard GMM asymptotic expansion (e.g. Andrews et al., 2020). [^7] The weighting matrix $W$ may be singular. The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints.
Here, $W$ may have columns of (effectively) zeros, which excludes the corresponding moments from the estimator. This occurs when researchers faces a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
The matrix $\Sigma$ depends on the experimental design (e.g., sample–size allocations across experimental moments). The constraints on $W$ and $\Sigma$ effectively limits the class of experiments, sample size allocations and potentially estimators the researcher may consider.
Assumption 3 (Constraint set).
Let $(W,\Sigma)\in\mathcal{D}^{\prime}$ for a set $\mathcal{D}^{\prime}$ so that $\Lambda^{\top}W\Lambda$ is invertible, $\Gamma_{\Lambda}(W)\Sigma\Gamma_{\Lambda}(W)^{\top}$ has uniformly bounded entries and is strictly positive definite.
Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments). This ensures that any subset of moments the researcher chooses to use is nondegenerate.
Example 5 (GMM moment selection).
Let $g(\theta)=(g^{\text{obs}}(\theta),\,g^{\text{exp}}_{2}(\theta),\,g^{\text{exp}}_{3}(\theta))^{\top}\in\mathbb{R}^{3}$, where the first component is observational and the other two are experimental. Suppose at most one experimental moment can be used. Selection is encoded by the GMM weight $W\in\mathbb{R}^{3\times 3}$. A simple example of class weighting matrices is
$W^{(2)}=\begin{pmatrix}w_{\mathrm{obs}}&\rho&0\\ \rho&w_{2}&0\\ 0&0&0\end{pmatrix}\qquad W^{(3)}=\begin{pmatrix}w_{\mathrm{obs}}&0&\rho\\ 0&0&0\\ \rho&0&w_{3}\end{pmatrix}$,
provided that $\Lambda^{\top}W\Lambda$ is invertible. ∎
Target and MSE.
Let the estimand be a linear functional $\tau(\theta)=\omega^{\top}\theta$ with $\omega\in\mathbb{R}^{d}$. Then
$\mathrm{MSE}_{b}(W,\Sigma)$ $=\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\;+\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega.$ (6)
The key distinction now is that the variance and bias regret depend on $\omega$ and $\Lambda$.
Definition 3 (Moment based variance and bias regret).
With a slight abuse of notation, let
$\alpha(W,\Sigma)\;\equiv\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega,\qquad\beta(W)\;\equiv\;\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}^{\,2}.$
and $\alpha^{\star},\beta^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$. The moment-based variance and bias regret are defined as $\alpha/\alpha^{\star},\beta/\beta^{\star}$ respectively.
The bias now depends on the interaction between the sensitivity $\omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased. The weighting matrix $W$ determines how these moments influence the downstream estimator $\hat{\theta}$.
As before, (and with a slight abuse of notation) we consider the adaptation regret of the following form:
$\mathcal{R}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\ \mathrm{MSE}_{b}(W^{\prime},\Sigma^{\prime})}.$ (7)
The next theorem generalizes Theorem 1 to this more general framework.
Theorem 2 (Regret characterization).
Let Assumptions 2, 3 hold, $\mathcal{R}$ be as defined in Equation (7),with $\alpha,\beta,\alpha^{\star},\beta^{\star}$ as in Definition 3. Consider an estimator as in Equation (5). Then
$\mathcal{R}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\beta(W)}{\beta^{\star}}\,\right\}.$ (8)
@@ 4 Generalization to moments selection
Theorem 2 shows how our framework generalizes to two related problems (i) choosing the weighting matrix $W$ for estimation and (ii) moment selection. The key insight is to show that the regret is quasi-convex under (approximate) linearity in the moments.